8.   Jimby has been taking care of Willy's garden each week for $25. Jimby's opportunity cost is $20, and Willy is willing to pay up to $35 to have someone take care of his garden each week. What is the maximum tax the government can impose on garden care without discouraging Jimby and Willy from continuing their mutually beneficial arrangement?
$  
Please enter a whole number, with no decimal point.
Answer #1

The maximum tax that the government can impose on garden care without discouraging Jimby and Willy from continuing their mutually beneficial arrangement is equal to

35-20 = 15
Answer #2

SOLUTION ;


John will not be at loss if he gets at least $20 (after tax)  equal to  opportunity cost. Willa is also not at a loss as he ays $35 only in any case.


Hence maximum tax that can be levied

= 35 - 20

= 15 ($) 


So maximum tax rate that can be levied on garden work 

= 15/35 * 100

= 42.86 %  (ANSWER)

answered by: Tulsiram Garg